OpenStudy (moonlitfate):

Use a right triangle to write the expression in algebraic form, given y = arccos x, 0 < y < pi/2 cot (y)

OpenStudy (anonymous):



OpenStudy (moonlitfate):

 That's the triangle that I have. I just haven't seen any of this in so long that I've forgotten.

OpenStudy (anonymous):

find the missing side using pythagoras you get \(\sqrt{1-x^2}\)

OpenStudy (anonymous):

then the cotangent is "adjacent over opposite" and you are done

OpenStudy (anonymous):

where you been?

OpenStudy (moonlitfate):

\[\frac{ x }{ \sqrt{1-x^2} }\]

OpenStudy (anonymous):

yes

OpenStudy (moonlitfate):

Thank you! Wow, I'm really rusty on all things math, but I guess there is where OPenStudy can help! :)

OpenStudy (anonymous):

rust is good nice to see you again

OpenStudy (moonlitfate):

Yeah, I will definitely be here. Calc. II is a beast...